OpenStudy (anonymous):

Name the algebraic property demonstrated in the example below. x + 5 = x + 5 Reflexive Property Symmetric Property Associative Property of Addition Distributive Property

OpenStudy (anonymous):

a = a (where a = x+5)

OpenStudy (anonymous):

It's communative property

OpenStudy (anonymous):

but there is no communative property?

OpenStudy (anonymous):

identity or reflexive, an item equals itself

OpenStudy (anonymous):

It's distributive, hope i helped :)

OpenStudy (anonymous):

identity is via the additive and multiplicative identities

OpenStudy (anonymous):

Distributive is eg Multiplication over addition

OpenStudy (anonymous):

Symmetric Property of Equality The following property: If if a = b then b = a. This is one of the equivalence properties of equality. Reflexive Property of Equality The property that a = a. One of the equivalence properties of equality.

OpenStudy (anonymous):

It's reflexive

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

So that's 3 votes for reflexive....

OpenStudy (anonymous):

aihgt guys thnks:)

OpenStudy (anonymous):

ASsociative isn't it, that's (1+2)+3 = 1+(2+3) Different

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@estudier: too bad math isn't solved by democracy!

OpenStudy (anonymous):

wait is it A?

OpenStudy (anonymous):

If A is the letter next to Reflexive property then it is A.

OpenStudy (anonymous):

yes tnk you:)